### v1

@@ 3.1 Fitting models
We analyzed each time–integrated spectrum with four different empirical functions: a smoothly broken power–law (SBPL), a double smoothly broken power–law (2SBPL, see Ravasio et al. 2018 for the description of their functional form), and their modified versions, SBPLCUTOFF and 2SBPLCUTOFF, respectively, which include the presence of a high–energy cutoff. Fig. 2 shows the four spectral shapes tested in this work, including the relative instrument we used to cover each energy range. The four fitting functions are described in the following.
@@
$highecut(E)=\begin{cases}1&\text{for $E\leq E_{\rm c}$}\\ e^{(E_{\rm c}-E)/E_{\rm fold}}&\text{for $E\geq E_{\rm c}$}\end{cases}$
@@
Figure 2: Representation of the four possible prompt emission spectral shapes and the corresponding spectral coverage of the GBM and the LAT instruments on board Fermi tested in this work, spanning 6 decades in energy. The typical GRBs spectrum is shown with the blue solid line, characterized by the usual $\alpha$, $\beta$ and $E_{\rm peak}$ parameters. Below the peak energy, the dashed black line shows the possible presence of a spectral break, modelled with the additional parameters $\alpha_{1}$, $\alpha_{2}$ and $E_{\rm break}$ (as shown in Ravasio et al. 2018, 2019). At higher energies, above $E_{\rm peak}$, the spectrum could show either the presence of a exponential cutoff, marked by $E_{\rm cutoff}$, or the extension of the power-law with slope $\beta$ (dashed black lines). The y-axis scale is arbitrary.
@@ 4 Results
In Table 2, we present the results of the time-integrated analysis from 10 keV up to 10 GeV of the selected sample of LLE-detected GRBs. For each GRB, the table lists the name, the best fit model (according to the $\Delta$AIC criterion reported above), the best fit parameters, the value of the PG-Statistic over the degrees of freedom, and, for those cases in which the spectrum is better fitted with the addition of a cutoff, the corresponding $\Delta$AIC. The reported errors are at the 1-$\sigma$ level.
@@
In the following sections we focus on the evidence of an exponential cutoff at high energies and on the slope of the spectrum above the peak energy.
@@ 4.1 High-energy spectral cutoff
In 9 out of 22 GRBs the addition of an exponential cutoff largely improves the fit, i.e. their time-integrated spectrum shows a statistically significant ($\Delta\rm AIC\geq 20.6$, corresponding to $\gtrsim 4\sigma$) exponential drop at high energies with respect to the basic model. The cutoff energy $E_{\rm cutoff}$ is well constrained in all the 9 spectra and its distribution spans a wide range of values, from $\sim$ 14 up to 298 MeV. In 4 GRB spectra the cutoff energy is above 100 MeV, while for the remaining 5 it falls in the LLE energy range (i.e. $E_{\rm cutoff}$ is below 100 MeV).
Fig. 3 shows an example of a GRB spectrum showing a cutoff at high–energy. In particular, the data refer to the $\nu F_{\nu}$ representation of the time-integrated spectrum of GRB 140206. The top plot shows the fit with the SBPL function (solid black line), which models the spectral shape with the following best–fit parameters: $\alpha=-1.01\pm 0.01$ , $\beta=-2.35\pm 0.01$ and $E_{\rm peak}=326.1_{-6.6}^{+8.8}$ keV. However, both the LLE and LAT data are not properly modelled by the high–energy power–law with slope $\beta$, as the power–law is overestimating the flux with respect to the data points from $\sim$ 80 MeV onwards (see the residual panel reported below the plots). The bottom plot shows the fit with the SBPLCUTOFF function. For comparison, the panel shows also the SBPL function without the exponential cutoff term (dashed black line). When fitted with the SBPLCUTOFF function, the spectral data are modelled by a larger peak energy $E_{\rm peak}=395.2_{-23.1}^{+29.1}$ keV and a harder $\beta$ parameter $\beta=-2.18\pm 0.03$ with respect to the fit with the SBPL. The exponential cutoff, parametrized by a cutoff energy at $E_{\rm cutoff}=88.3^{+18.2}_{-12.5}$ MeV, allows to properly model the high-energy data and absorbs the previously visible trend in the residuals. The addition of an exponential cutoff provides a highly significant improvement of the fit, statistically evaluated in this case by a $\Delta$AIC = 69.1.
Figure 3: Example of a time-integrated spectrum showing a cutoff at high–energy. Data correspond to the signal detected by different instruments (colour-coded as shown in the legend) for GRB 140206. The top plot shows the fit with the SBPL function, represented by the solid black line. The bottom plot shows the same spectral data fitted with the addition of an exponential cutoff, namely with the SBPLCUTOFF function, represented with the solid black line. The dashed black line displays the SBPL function without the exponential cutoff term. The parameters of the best–fitting models and their 1-$\sigma$ errors are reported in each plot (see also Tab. 2). The bottom panel of each plot shows the residuals of the spectral fit (computed as the distance of the data from the model in units of the errors). Data have been rebinned for graphical purposes.
@@ 4.2 Beta distribution
For all the models tested, the shape of the spectrum above the peak energy is modelled by a power–law with slope $\beta$. This allows us to study the high–energy spectral slope for each GRB of our sample, regardless the presence of a cutoff.
By combining the results from all the best–fitting models shown in Table 2, we built the distribution of the $\beta$ parameter, which we reported with the black empty histogram in Fig. 4. Note the distribution does not show values greater than $-2$. This is simply due to the hard limit we set on the $\beta$ parameter, in order to have a peak in the $\nu F_{\nu}$ spectrum. All the spectra analyzed in our sample showed a well defined peak in the broad energy band studied, such that, if the constraint on the $\beta$ parameter would be relaxed, we would still have $\beta$ values smaller than $-2$.
The distribution of $\beta$ shown in Fig. 4 spans a wide range, from -4.60 to -2.06, and it is has a median value of $\langle\beta\rangle$ = -2.39 (standard deviation of 0.56). The average error on the $\beta$ values together with the median value are represented by the black horizontal bar and the point over-plotted above the histograms.
Figure 4: Distributions of the high-energy slope $\beta$ (photon index) for the 22 GRBs spectra analyzed. The different best–fitting models are represented by different colors (stacked on top of each other), while the black empty histogram represents the total distribution of $\beta$ values derived from the spectral fits performed in this work. On the top of the distributions it is over–plotted the median value and the mean error on the $\beta$ parameter for the whole sample. For comparison, the distribution of the $\beta$ values derived from the Fermi/GBM Catalog is represented by red dashed histogram. The distributions have been normalized to the respective sample size.
@@
We now compare the distributions of $\beta$ for the 13 GRB spectra not showing a cutoff (yellow and blue histograms in Fig. 4) with the 9 GRBs showing a cutoff (orange and green distributions). The mean value of $\beta$ for the 13 GRB spectra not showing a cutoff is $\langle\beta_{\rm no\,cutoff}\rangle=-2.81$ ($\sigma$= 0.64), while for the 9 GRBs showing a cutoff is $\langle\beta_{\rm cutoff}\rangle=-2.33$ ($\sigma$= 0.19). The small difference between the two distributions could be explained as an effect due to the presence of the exponential cutoff. In fact, under the assumption that the intrinsic high–energy slope is the same, when an exponential cutoff is present in the spectrum, the power–law segment between the peak energy and the cutoff energy is truncated earlier, allowing to constrain $\beta$ on a smaller energy range. This effect, combined with the broad curvature of the spectrum, likely provides shallower (i.e. harder) values of the high–energy slope of GRBs showing a cutoff.
@@
The diversity in the $\beta$ distributions could be ascribed to the addition of the LLE and LAT data to the GBM ones in our analysis. While in the Fermi/GBM Catalog are reported the parameters of the fit obtained considering spectral data up to 40 MeV, in this analysis we extended the characterization of the spectra up to 10 GeV. This may suggest that high-energy slope of the GRBs prompt emission spectrum is indeed softer than what commonly deduced from fits made with lower energy data alone (e.g. from BATSE or GBM). A similar conclusion has been drawn also by Fermi Large Area Telescope Team et al. 2012, where it was shown that in 24 bursts co-detected by GBM and LAT the $\beta$ values are systematically softer than the values found from fitting the GBM data alone. Also, when performing the joint fit of GBM and LAT data, they found no cases of spectra with $\beta>-2$, in agreement with the constraint we set on this parameter. Perhaps less dominant, but still of impact on our $\beta$ distribution, could be also the effect of the use of more complex fitting functions in our analysis, as compared to the simpler Band and SBPL function used in the Catalog.
@@ 5 Interpretation
In this section we provide the interpretation of the results obtained from the analysis of the high-energy extension of 22 GRBs bright prompt emission spectra, described in the previous section.
This section is divided in two parts. The first part regards the evidence of the exponential cutoff, which will be interpreted as the sign of pair-production opacity, allowing us to derive the bulk Lorentz factor of the jet. The alternative interpretation, involving the maximum photon energy for synchrotron emission, will be discussed in Sec. 6. The second part of this section concerns the results on the slope $\beta$ of the high-energy power–law, which allows to put valuable constrain on the slope of the underlying distribution of the accelerated particles.
@@ 5.1 Estimate of $\Gamma$ from pair-production opacity
By interpreting the observed exponential cutoff as due to the absorption caused by the photon–photon process, producing electron–positron pairs, we can infer an estimate of the bulk Lorentz factor $\Gamma$ of the jet during the prompt emission phase (e.g. Piran 1999; Lithwick & Sari 2001).
@@
Eq. 1 depends mainly on the number density at threshold and weakly on the spectral shape of the number photon density. This is because of the steep decrease of the cross section $\sigma_{\gamma\gamma}(\epsilon_{T}^{\prime})$ with energy. Furthermore, for typical spectral indices $\beta_{e}$, also $n^{\prime}(\epsilon_{T}^{\prime})$ decreases with energy. Svensson 1987 numerically calculated the correct $\tau_{\gamma\gamma}$ encapsulating the spectral index dependence into the function $\eta(\beta_{e})$ which can be approximated with (see his Eq. B5):
$\eta(\beta_{e})={7\over 6}\,{1\over(2+\beta_{e})(1+\beta_{e})^{5/3}}$ (2)
@@
For the computation of the bulk Lorentz factor, we need the redshift $z$ of the source, which is required to compute the isotropic luminosity, the rest frame energies and the variability timescale in the equation. For 9 GRBs in our sample, the redshift has been spectroscopically measured (see Table 1), and we used it in the computation of $\Gamma$. For those GRBs without a redshift measurement, we used $z=2$ as representative value of the redshift distribution of long GRBs [^11] We repeated the analysis assuming 4 different redshift values, i.e. $z$in [0.5, 1 , 3, 4], finding that the corresponding values of $\Gamma$ are on average a factor [0.50, 0.70, 1.25, 1.48] times the ones derived assuming $z=2$. The choice of not testing redshift values higher than $z=4$ is motivated by the extreme values of the corresponding $E_{\rm iso}(\gtrsim 4\times 10^{55}$ erg) otherwise implied by higher redshift for the considerably bright GRBs selected in this work.. Each GRBs lightcurve shows different temporal behaviours and varies on different timescales. For the sake of simplicity, we assume the same value of the variability timescale for all the GRBs, namely $t_{\rm var}$ = 0.1 s, which corresponds to the mean value of the variability timescales for long bursts observed by Fermi (Golkhou et al. 2015).
@@
Setting $\tau_{\gamma\gamma}(E_{\rm cutoff})=1$ in Eq. 6 allows us to derive an estimate of $\Gamma$ for those 9 GRBs showing a cutoff in their spectrum, as reported in Sec. 4. For the remaining 13 GRBs not showing a cutoff, we used the lower–limit on $E_{\rm cutoff}$ derived from the spectral fits in order to set a lower–limit on their $\Gamma$. The uncertainties on the measured quantities $E_{\rm cutoff}$, $E_{\rm peak}$, $L_{\rm iso}$, $\alpha_{\rm e}$ and $\beta_{\rm e}$ give a relative uncertainty of the order of $\sim 4$% on the estimates of $\Gamma$. These values of $\Gamma$ must be compared with $\Gamma_{\rm max}=(1+z)\,E_{\rm cutoff}/m_{e}c^{2}$. This limit corresponds to the maximum bulk Lorentz factor attainable for a given observed cutoff energy, since the cutoff energy in the comoving frame would be at the self-annihilation threshold ($E_{\rm cutoff}^{\prime}=m_{e}c^{2}$)  [^12] This is due the fact that, when we observe a cutoff energy in the spectrum, it could correspond at least to $m_{e}c^{2}$ in the comoving frame, which directly gives a limit to the maximum value of bulk Lorentz factor $\Gamma$.. The bulk Lorentz factor is then the minimum of the obtained values. We note that, for all the 9 GRBs showing a cutoff, the estimate of $\Gamma$ is below the computed maximum value $\Gamma_{\rm max}$: we therefore consider the value derived from Eq. 6 as the estimate of their bulk Lorentz factor.
@@
Figure 5: Distribution of bulk Lorentz factor $\Gamma$ obtained from the spectral analysis of the 22 bright GRB co-detected by GBM and LAT considered in this work. The values of $\Gamma$ are derived by interpreting the high–energy cutoff (either a measurement of or a lower–limit of) as a sign of pair-production opacity, following Eq. 7. The blue filled histogram represents the measurement of $\Gamma$ for the 9 GRBs of the sample showing evidence of high–energy cutoff in their prompt emission spectra. The dashed red histogram (with rightward arrows) represents the lower–limit on $\Gamma$ from the 13 GRBs not showing a cutoff in their spectrum (see the text for details).
@@
Table 3 shows the values and the lower–limits of $\Gamma$ derived from Eq. 7 for each GRB in our sample, either with or without the detection of an exponential cutoff in the spectrum. The distribution of the estimates of $\Gamma$ for the 9 GRBs showing a cutoff is shown with the blue filled histogram in Fig. 5, together with the distribution of the lower–limits on $\Gamma$ for the 13 GRBs not showing a cutoff (red dashed histogram). The distribution of measured $\Gamma$ values is centered around $\langle\Gamma\rangle=229\,(\sigma=68)$, while the distribution of the lower–limits on $\Gamma$ has a mean value of $\langle\Gamma\rangle=317\,(\sigma=157)$. The main reason for the difference in the two distributions can be due to the fact that, for the 13 GRBs without the detection of a cutoff, it has been used the lower–limit on the cutoff energy to derive $\Gamma$ and their values are of the order of several hundreds of MeV (typically a factor $\sim$ 5–6 times larger than the estimates of $E_{\rm cutoff}$ for the 9 GRBs with a detected exponential cutoff). The lower–limits on $E_{\rm cutoff}$ imply that there could be a cutoff at higher energies. Speculating on the presence of such high-energy cutoffs, which would not be observable with the current instruments, this would in turn imply that the distribution of $\Gamma$ would be broader, extending towards larger values. This is indeed compatible with what inferred from the afterglow observations, as it will be discussed in Section 6.3.

### v2

@@ 3.2 Fitting models
We analyzed each time–integrated spectrum with four different empirical functions: a smoothly broken power law (SBPL), a double smoothly broken power law (2SBPL, see Ravasio et al. 2018 for the description of their functional form), and their modified versions, SBPLCUTOFF and 2SBPLCUTOFF, respectively, which include the presence of a high–energy cutoff. Fig. 2 shows the four spectral shapes tested in this work, including the relative instrument we used to cover each energy range. The four fitting functions are described in the following.
@@
highecut(E)={1&for EEc
e(EcE)/Efoldfor EEc
[…]
@@
Figure 2: Representation of the four possible prompt emission spectral shapes and the corresponding spectral coverage of the GBM and the LAT instruments on board Fermi tested in this work, spanning 6 decades in energy. The typical GRBs spectrum is shown with the blue solid line, characterized by the usual $\alpha$, $\beta$ and $E_{\rm peak}$ parameters. Below the peak energy, the dashed black line shows the possible presence of a spectral break, modeled with the additional parameters $\alpha_{1}$, $\alpha_{2}$ and $E_{\rm break}$ (as shown in Ravasio et al. 2018, 2019). At higher energies, above $E_{\rm peak}$, the spectrum could show either the presence of an exponential cutoff, marked by $E_{\rm cutoff}$, or the extension of the power-law with slope $\beta$ (dashed black lines). The y-axis scale is arbitrary.
@@ 4 Results
In Table 2, we present the results of the time-integrated analysis from 10 keV up to 10 GeV of the selected sample of LLE-detected GRBs. For each GRB, the table lists the name, the best fit model (according to the $\Delta$AIC criterion reported above), the best fit parameters, the value of the PG-Statistic over the degrees of freedom, and, for those cases in which the spectrum is better fitted with the addition of a cutoff, the corresponding $\Delta$AIC. The reported errors are at the 1-$\sigma$ level. In the following sections, we focus on the evidence of an exponential cutoff at high energies and on the slope of the spectrum above the peak energy.
@@ 4.1 High-energy spectral cutoff
In nine out of 22 GRBs the addition of an exponential cutoff largely improves the fit, that is, their time-integrated spectrum shows a statistically significant ($\Delta\rm AIC\geq 20.6$, corresponding to $\gtrsim 4\sigma$) exponential drop at high energies with respect to the basic model. The cutoff energy $E_{\rm cutoff}$ is well constrained in all the 9 spectra and its distribution spans a wide range of values, from $\sim$ 14 up to 298 MeV. In 4 GRB spectra, the cutoff energy is above 100 MeV, while for the remaining 5 it falls in the LLE energy range (i.e., $E_{\rm cutoff}$ is below 100 MeV).
Fig. 3 shows an example of a GRB spectrum showing a cutoff at high–energy. In particular, the data refer to the $\nu F_{\nu}$ representation of the time-integrated spectrum of GRB 140206. The top plot shows the fit with the SBPL function (solid black line), which models the spectral shape with the following best–fit parameters: $\alpha=-1.01\pm 0.01$ , $\beta=-2.35\pm 0.01$ and $E_{\rm peak}=326.1_{-6.6}^{+8.8}$ keV. However, both the LLE and LAT data are not properly modeled by the high–energy power–law with slope $\beta$, as the power–law overestimates the flux with respect to the data points from $\sim$ 80 MeV onward (see the residual panel reported below the plots). The bottom plot shows the fit with the SBPLCUTOFF function. For comparison, the panel shows also the SBPL function without the exponential cutoff term (dashed black line). When fitted with the SBPLCUTOFF function, the spectral data are modeled by a larger peak energy $E_{\rm peak}=395.2_{-23.1}^{+29.1}$ keV and a harder $\beta$ parameter $\beta=-2.18\pm 0.03$ with respect to the fit with the SBPL. The exponential cutoff, parametrized by a cutoff energy at $E_{\rm cutoff}=88.3^{+18.2}_{-12.5}$ MeV, allows to properly model the high-energy data and absorbs the previously visible trend in the residuals. The addition of an exponential cutoff provides a highly significant improvement of the fit, statistically evaluated in this case by a $\Delta$AIC = 69.1.
Figure 3: Example of a time-integrated spectrum showing a cutoff at high energy. Data correspond to the signal detected by different instruments (color-coded as shown in the legend) for GRB 140206. The top plot shows the fit with the SBPL function, represented by the solid black line. The bottom plot shows the same spectral data fitted with the addition of an exponential cutoff, namely with the SBPLCUTOFF function, represented with the solid black line. The dashed black line displays the SBPL function without the exponential cutoff term. The parameters of the best–fitting models and their 1-$\sigma$ errors are reported in each plot (see also Tab. 2). The bottom panel of each plot shows the residuals of the spectral fit (computed as the distance of the data from the model in units of the errors). Data have been rebinned for graphical purposes.
@@ 4.2 Beta distribution
For all the models tested, the shape of the spectrum above the peak energy is modeled by a power–law with slope $\beta$. This allows us to study the high–energy spectral slope for each GRB of our sample, regardless of the presence of a cutoff.
By combining the results from all the best–fitting models shown in Table 2, we built the distribution of the $\beta$ parameter, which we reported with the black empty histogram in Fig. 4. It is important to note that the distribution does not show values greater than $-2$. This is simply due to the hard limit we set on the $\beta$ parameter, in order to have a peak in the $\nu F_{\nu}$ spectrum. All the spectra analyzed in our sample showed a well-defined peak in the broad energy band studied, such that, if the constraint on the $\beta$ parameter would be relaxed, we would still have $\beta$ values smaller than $-2$.
The distribution of $\beta$ shown in Fig. 4 spans a wide range, from -4.60 to -2.06, and it is has a median value of $\langle\beta\rangle$ = -2.39 (standard deviation of 0.56). The average error on the $\beta$ values together with the median value of the distribution are represented by the black horizontal bar and the point over-plotted above the histograms.
Figure 4: Distributions of the high-energy slope $\beta$ (photon index) for the 22 GRBs’ spectra analyzed. The different best–fitting models are represented by different colors (stacked on top of each other), while the black empty histogram represents the total distribution of $\beta$ values derived from the spectral fits performed in this work. On the top of the distributions it is over–plotted the median value and the mean error on the $\beta$ parameter for the whole sample. For comparison, the distribution of the $\beta$ values derived from the Fermi/GBM catalog is represented by the red dashed histogram. The distributions have been normalized to the respective sample size.
@@
We now compare the distributions of $\beta$ for the 13 GRB spectra not showing a cutoff (yellow and blue histograms in Fig. 4) with the 9 GRBs showing a cutoff (orange and green distributions). The median value of $\beta$ for the 13 GRB spectra not showing a cutoff is $\langle\beta_{\rm no\,cutoff}\rangle=-2.69$ ($\sigma$= 0.64), while for the 9 GRBs showing a cutoff is $\langle\beta_{\rm cutoff}\rangle=-2.33$ ($\sigma$= 0.19). The small difference between the two distributions could be explained as an effect due to the presence of the exponential cutoff. In fact, under the assumption that the intrinsic high–energy slope is the same, when an exponential cutoff is present in the spectrum, the power–law segment between the peak energy and the cutoff energy is truncated earlier, allowing to constrain $\beta$ on a smaller energy range. This effect, combined with the broad curvature of the spectrum, likely provides shallower (i.e., harder) values of the high–energy slope of GRBs showing a cutoff.
@@
The diversity in the $\beta$ distributions could be ascribed to the addition of the LLE and LAT data to the GBM ones in our analysis. While in the Fermi/GBM catalog are reported the parameters of the fit obtained considering spectral data up to 40 MeV, in this analysis we extended the characterization of the spectra up to 10 GeV. This may suggest that high-energy slope of the GRBs prompt emission spectrum is indeed softer than what is commonly deduced from fits made with lower energy data alone (e.g., from BATSE or GBM). A similar conclusion has been drawn also by Fermi Large Area Telescope Team et al. 2012, where it was shown that in 24 bursts co-detected by GBM and LAT the $\beta$ values are systematically softer than the values found from fitting the GBM data alone. Also, when performing the joint fit of GBM and LAT data, they found no cases of spectra with $\beta>-2$, in agreement with the constraint we set on this parameter. Perhaps less dominant, but still of impact on our $\beta$ distribution, could be also the effect of the use of more complex fitting functions in our analysis, as compared to the simpler Band and SBPL function used in the catalog.
@@ 5 Interpretation
In this section, we provide the interpretation of the results obtained from the analysis of the high-energy extension of 22 GRBs bright prompt emission spectra, described in the previous section. This section is divided into two parts. The first part regards the evidence of the exponential cutoff, which will be interpreted as the sign of pair-production opacity, allowing us to derive the bulk Lorentz factor of the jet. The alternative interpretation, involving the maximum photon energy for synchrotron emission, will be discussed in Sec. 6. The second part of this section concerns the results on the slope $\beta$ of the high-energy power–law, which allows to put valuable constrain on the slope of the underlying distribution of the accelerated particles.
@@ 5.1 Estimate of $\Gamma$ from pair-production opacity
By interpreting the observed exponential cutoff as due to the absorption caused by the photon–photon process, producing electron–positron pairs, we can infer an estimate of the bulk Lorentz factor $\Gamma$ of the jet during the prompt emission phase (e.g., Piran 1999; Lithwick & Sari 2001).
@@
Eq. 1 mainly depends on the number density at threshold and weakly on the spectral shape of the number photon density. This is because of the steep decrease of the cross-section $\sigma_{\gamma\gamma}(\epsilon_{T}^{\prime})$ with energy. Furthermore, for typical spectral indices $\beta_{e}$, also $n^{\prime}(\epsilon_{T}^{\prime})$ decreases with energy. Svensson 1987 numerically calculated the correct $\tau_{\gamma\gamma}$ encapsulating the spectral index dependence into the function $\eta(\beta_{e})$ which can be approximated with (see his Eq. B5) the following:
$\eta(\beta_{e})={7\over 6}\,{1\over(2+\beta_{e})(1+\beta_{e})^{5/3}}.$ (2)
@@
For the computation of the bulk Lorentz factor, we need the redshift $z$ of the source, which is required to compute the isotropic luminosity, the rest frame energies and the variability timescale in the equation. For 9 GRBs in our sample, the redshift has been spectroscopically measured (see Table 1), and we used it in the computation of $\Gamma$. For those GRBs without a redshift measurement, we used $z=2$ as representative value of the redshift distribution of long GRBs [^11] We repeated the analysis assuming four different redshift values, i.e., $z$in [0.5, 1 , 3, 4], finding that the corresponding values of $\Gamma$ are on average a factor [0.50, 0.70, 1.25, 1.48] times the ones derived assuming $z=2$. The choice of not testing redshift values higher than $z=4$ is motivated by the extreme values of the corresponding $E_{\rm iso}(\gtrsim 4\times 10^{55}$ erg) otherwise implied by higher redshift for the considerably bright GRBs selected in this work.. Each GRBs light curve shows different temporal behaviors and varies on different timescales. For the sake of simplicity, we assume the same value of the variability timescale for all the GRBs, namely $t_{\rm var}$ = 0.1 s, which corresponds to the mean value of the variability timescales for long bursts observed by Fermi (Golkhou et al. 2015).
@@
Setting $\tau_{\gamma\gamma}(E_{\rm cutoff})=1$ in Eq. 6 allows us to derive an estimate of $\Gamma$ for those 9 GRBs showing a cutoff in their spectrum, as reported in Sec. 4. For the remaining 13 GRBs not showing a cutoff, we used the lower–limit on $E_{\rm cutoff}$ derived from the spectral fits in order to set a lower–limit on their $\Gamma$. The uncertainties on the measured quantities $E_{\rm cutoff}$, $E_{\rm peak}$, $L_{\rm iso}$, $\alpha_{\rm e}$ and $\beta_{\rm e}$ give a relative uncertainty on the order of $\sim 4$% on the estimates of $\Gamma$. These values of $\Gamma$ must be compared with $\Gamma_{\rm max}=(1+z)\,E_{\rm cutoff}/m_{e}c^{2}$. This limit corresponds to the maximum bulk Lorentz factor attainable for a given observed cutoff energy, since the cutoff energy in the comoving frame would be at the self-annihilation threshold ($E_{\rm cutoff}^{\prime}=m_{e}c^{2}$)  [^12] This is due the fact that, when we observe a cutoff energy in the spectrum, it could correspond at least to $m_{e}c^{2}$ in the comoving frame, which directly gives a limit to the maximum value of bulk Lorentz factor $\Gamma$.. The bulk Lorentz factor is then the minimum of the obtained values. We note that, for all the 9 GRBs showing a cutoff, the estimate of $\Gamma$ is below the computed maximum value $\Gamma_{\rm max}$: we therefore consider the value derived from Eq. 6 as the estimate of their bulk Lorentz factor.
@@
Figure 5: Distribution of bulk Lorentz factor $\Gamma$ obtained from the spectral analysis of the 22 bright GRBs that were co-detected by GBM and LAT and considered for this work. The values of $\Gamma$ are derived by interpreting the high–energy cutoff (either a measurement of or a lower–limit of) as a sign of pair-production opacity, following Eq. 7. The blue-filled histogram represents the measurement of $\Gamma$ for the 9 GRBs of the sample showing evidence of high–energy cutoff in their prompt emission spectra. The dashed red histogram (with rightward arrows) represents the lower–limit on $\Gamma$ from the 13 GRBs not showing a cutoff in their spectrum (see the text for details).
@@
Table 3 shows the values and the lower–limits of $\Gamma$ derived from Eq. 7 for each GRB in our sample, either with or without the detection of an exponential cutoff in the spectrum. The distribution of the estimates of $\Gamma$ for the 9 GRBs showing a cutoff is shown with the blue filled histogram in Fig. 5, together with the distribution of the lower–limits on $\Gamma$ for the 13 GRBs not showing a cutoff (red dashed histogram). The distribution of measured $\Gamma$ values is centered around $\langle\Gamma\rangle=229\,(\sigma=68)$, while the distribution of the lower–limits on $\Gamma$ has a mean value of $\langle\Gamma\rangle=317\,(\sigma=157)$. The main reason for the difference in the two distributions can be due to the fact that, for the 13 GRBs without the detection of a cutoff, it has been used the lower–limit on the cutoff energy to derive $\Gamma$ and their values are on the order of several hundreds of MeV (typically a factor $\sim$ 5–6 times larger than the estimates of $E_{\rm cutoff}$ for the 9 GRBs with a detected exponential cutoff). The lower–limits on $E_{\rm cutoff}$ imply that there could be a cutoff at higher energies. Speculating on the presence of such high-energy cutoffs, which would not be observable with the current instruments, this would in turn imply that the distribution of $\Gamma$ would be broader, extending toward larger values. This is indeed compatible with what inferred from the afterglow observations, as it will be discussed in Section 6.3.